1. 500 mL C2H6 was measured at STP, Determine how many grams of C2H6 were measured.

2. The mass percentage composition of sodium metal in 58.44 g NaCl is about 39.34%, Determine the percentage composition of sodium in 1000 g of NaCl.
1
Expert's answer
2020-08-31T12:47:57-0400

Question 1

"\nu=\frac{m}{M}=\frac{V}{V_m}" whence "m=\frac{VM}{V_m}"

"M(C_2H_6)=110\frac{gram}{mole}"

Molar volume "V_m=22.4" (L)

"m=\frac{500*10^{-3}*110}{22.4}=2.5" (g)


Question 2:

If W is the mass percentage, it has the linear dependence from mass.

"W_1-m_1\nW_2-m_2"

"\frac{W_1}{W_2}=\frac{m_1}{m_2}"

"W_2=\frac{m_2}{m_1}W_1"

"W_2=\frac{1000}{58.44}*0.3934=0.023=6.73" (it's a very big number)